OpenStudy (anonymous):

Can someone explain inequalities to me.. y² - 8y - 9 ≤ 0 (y + 1)(y - 9) ≤ 0 but why is the answer -1 ≤ y ≤ 9 and not y ≤ 9, y ≤ -1 ?

ganeshie8 (ganeshie8):

(y + 1)(y - 9) ≤ 0 First, we notice that product of two number is negative ONLY when both the numbers are of opposite signs.

ganeshie8 (ganeshie8):

for example : (2)(-3) = -6 the product is negative cuz `2` and `-3` are of opposite signs.

OpenStudy (anonymous):

uhm.. i think that doesnt answer my question but thanks anyway! :D

ganeshie8 (ganeshie8):

thats just a starter :) (y + 1)(y - 9) ≤ 0 can be broken into two cases : 1) y+1 <= 0 and y-9 >= 0 2) y+1 >=0 and y-9 <= 0

ganeshie8 (ganeshie8):

solve each case and see what you get

OpenStudy (anonymous):

uhm.. i got 1. y <= -1 and y >= 9 2. y >= -1 and y <= 9

ganeshie8 (ganeshie8):

Excellent ! take a look at 1st case, the solution we have gotten is just nonsense (why ?) so discard it only the 2nd case is valid.

OpenStudy (anonymous):

i noticed that in each case there are ≤ and ≥. can i ask how does the ≥ come about? Since the original equation only gave ≤, where (y + 1)(y - 9) ≤ 0

ganeshie8 (ganeshie8):

thats a good question :)

ganeshie8 (ganeshie8):

actually there is another method which won't make u ask these questions, but i don't like that method... lets look at your question :)

ganeshie8 (ganeshie8):

we have below inequality to solve : (y + 1)(y - 9) ≤ 0

ganeshie8 (ganeshie8):

It is a product of two individual factors: (y+1) and (y-9) what do we know about the `signs` of these individual factors ?

ganeshie8 (ganeshie8):

For the product to be negative, these factors must have opposite `signs` right ? (if one factor is <=0, then other factor has to be >=0)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(y+1) meaning that the answer will be -1 and 9 for (y-9)

ganeshie8 (ganeshie8):

yes so did u get how we got these two cases ? (y + 1)(y - 9) ≤ 0 1) y+1 <= 0 and y-9 >= 0 2) y+1 >=0 and y-9 <= 0

ganeshie8 (ganeshie8):

if it helps, we're using this fact : \(\large - \times + = -\) \(\large + \times - = -\)

OpenStudy (anonymous):

waaat :O

OpenStudy (anonymous):

nvm im too dumb.. cuz all i see is to shift from left side of the equation to right side of the equation ._.

OpenStudy (anonymous):

thats why i cant see how the 2 cases and the ≥ come about @@

OpenStudy (anonymous):

can you try telling me the other method maybe i wont see how does the ≥ come about ? :x

ganeshie8 (ganeshie8):

Ahh I can see how that could be confusing...

ganeshie8 (ganeshie8):

but i don't really like the other method as it involves more new terms and more explanation - IMHO this method is simple but it has to makes sense first -.-

OpenStudy (anonymous):

oh okie >.<

OpenStudy (mathstudent55):

@HatcrewS Let me try to explain to you what you don't understand. @ganeshie8 did a great job with this problem, and you followed it well. Your question above is: "i noticed that in each case there are ≤ and ≥. can i ask how does the ≥ come about? Since the original equation only gave ≤, where (y + 1)(y - 9) ≤ 0"

OpenStudy (mathstudent55):

When you have the product (y + 1)(y - 9) ≤ 0, think of the "≤ 0" part as meaning it's 0 or a negative number. A number less than zero is negative.

OpenStudy (mathstudent55):

The question is what two numbers can you multiply together and get a negative result? The answer is one must be positive and one must be negative. When you get an answer in the form y >= -1 and y <= 9, instead of thinking of the inequality sign having flipped, think of it as meaning the answer is any number between -1 and 9. It's just that to write that down we need to flip the inequality sign.

OpenStudy (anonymous):

omg #mind blown. thank you so much @mathstudent55! i have a much better understanding now :)

OpenStudy (mathstudent55):

You're welcome.